# DragPoser: Motion Reconstruction from Variable Sparse Tracking Signals via Latent Space Optimization

### v1

@@ Abstract.
High-quality motion reconstruction that smoothly and accurately follows the user’s movements can be achieved by high-end mocap systems consisting of many sensors. However, obtaining such animation quality with fewer input devices is gaining popularity as it brings mocap closer to the general public. The main challenges with using fewer trackers include the loss of end-effector accuracy in learning-based approaches, or the lack of naturalness and smoothness in IK-based solutions. In addition, such systems are often finely tuned to a specific number of trackers and are highly sensitive to missing data e.g., in scenarios where a sensor is occluded or malfunctions. In response to these challenges, we introduce DragPoser, a novel deep-learning-based motion reconstruction system that accurately represents hard and dynamic on-the-fly constraints, attaining real-time high end-effectors position accuracy. This is achieved through a pose optimization process within a structured latent space. Our system requires only one-time training on a large human motion dataset, and then constraints can be dynamically defined as losses, while the pose is iteratively refined by computing the gradients of these losses within the latent space. To further enhance our approach, we incorporate a Temporal Predictor network, which employs a Transformer architecture to directly encode temporality within the latent space. This network ensures the pose optimization is confined to the manifold of valid poses and also leverages past pose data to predict temporally coherent poses. Results demonstrate that DragPoser surpasses both IK-based and the latest data-driven methods in achieving precise end-effector positioning, while it produces natural poses and temporally coherent motion. In addition, our system showcases robustness against on-the-fly constraint modifications, and exhibits exceptional adaptability to various input configurations and changes, even in scenarios where data for specific end-effectors are completely absent. The versatility and efficiency of DragPoser, yielding high-quality results, not only extend its accessibility to a diverse user base but also enhance its performance in real-world applications.
Motion Tracking, Character Animation, Wearable Devices, Sparse Data
submissionid: 878copyright: noneccs: Computing methodologies Motion captureccs: Computing methodologies Motion processingccs: Computing methodologies Animationccs: Computing methodologies Learning paradigms
Figure 1. Our approach reconstructs complete full-body poses using only a few sensors (highlighted in green). In contrast to state-of-the-art learning-based methods, our technique achieves animation smoothness while at the same time incorporating precise end-effector position constraints (left). Furthermore, it offers the flexibility to seamlessly incorporate additional real-time constraints, enabling graceful handling of scenarios such as missing sensor data (middle). Notably, our versatile one-for-all architecture has the capability to adapt to arbitrary combinations of sensors (right), eliminating the need for modifications to the network’s architecture or retraining.
1. Introduction
@@
Human motion reconstruction and editing have gained significant attention in diverse industries such as entertainment, sports, and rehabilitation, with applications spanning virtual reality (VR), augmented reality (AR), training, education, simulations, and collaborative work. In professional settings, motion reconstruction typically relies on high-quality optical motion capture (mocap) systems or Inertial Measurement Unit (IMU)-based systems equipped with numerous sensors. These systems, although effective, come with significant costs and necessitate extensive calibrations and technical expertise. However, recent trends indicate a growing demand for cost-effective and minimally intrusive mocap systems that leverage consumer-grade hardware for motion reconstruction. This shift in focus aims to cater to applications in the Metaverse or home users, targeting affordability and accessibility. Conversely, motion editing often involves the use of non-learned inverse kinematics (IK) methods such as CCD (Kenwright, 2012) or FABRIK (Aristidou and Lasenby, 2011). While IK enables the rapid creation of animations using end-effectors, it frequently leads to time-incoherent or unnatural human motion due to the inherently underdetermined nature of the problem.
@@
With the growing availability of extensive repositories of motion capture data and the rapid advancements in deep learning, various data-driven methodologies have emerged to grapple with the intricacies and challenges of human motion reconstruction. Recent studies have focused on the reconstruction of poses from sparse input (Ponton et al., 2023; Yi et al., 2023; Winkler et al., 2022), aiming to reduce reliance on expensive motion capture suits. These methodologies, leveraging their learning from vast and precise mocap datasets, reconstruct human motion that is characterized by both temporal coherence and naturalness, overcoming the shortcomings of traditional IK (Oreshkin et al., 2021; Agrawal et al., 2023). However, they often encounter limitations arising from their rigid neural network architectures, restricting users to a specific set of sparse inputs or constraints. Additionally, they are sensitive to faulty sensors or tracking losses. Despite their capacity to reconstruct smooth animations, they commonly face challenges in dealing with hard constraints, often requiring the fallback to traditional IK methods for refining final poses (Jiang et al., 2022a; Ponton et al., 2022).
@@
This paper introduces DragPoser, an innovative deep-learning-based system designed for motion reconstruction. Unlike conventional methods that rely on direct full-body pose prediction from sparse inputs, our method leverages a structured latent space and employs a pose optimization process to accurately represent both hard and dynamic constraints. The key difference lies in our utilization of an optimization process within the learned latent space, enabling dynamic guidance in the pose search. The system undergoes a singular training phase on an extensive human motion dataset. Subsequently, constraints are dynamically defined as losses, and the pose is iteratively refined by computing gradients within the latent space. To ensure temporal coherence and constrain pose optimization within the valid pose manifold, we integrate a Temporal Predictor network, which directly encodes temporality into the latent space. This departure from traditional direct pose prediction methods enables a more versatile and dynamic exploration of poses, providing increased generality and improved accuracy when enforcing rigid constraints on end effector positions.
@@
We extensively tested and compared our method against state-of-the-art methods in a variety of challenging motion sequences from public datasets, with different sensor configurations. Our results show that DragPoser outperforms both traditional IK-based methods and the newest data-driven approaches in nearly all error metrics. DragPoser excels at providing accurate end-effector positioning (as in traditional IK systems) but at the same time, it generates natural poses and temporally coherent motion, a feat that data-driven methods typically achieve only in the absence of hard constraints. DragPoser, however, boasts another distinct advantage: its adaptability to hardware changes. We showcase this adaptability by demonstrating DragPoser’s robust capability to effectively handle missing input data caused by occlusions, faulty sensors, or communication problems. Additionally, users can dynamically add constraints in real-time to target specific types of motions. This dual adaptability not only reinforces the system’s resilience but also enhances its performance in real-world applications. Upon acceptance, the code will be released to facilitate ongoing development, comparisons, and evaluations.
2. Related Work
@@
2.1. Motion Capture with Sparse Input
@@
Recent research in human motion reconstruction has focused on utilizing a reduced number of Inertial Measurement Units (IMUs) attached to the body. This approach eliminates the need for external sensors or cameras, requires no line-of-sight, and operates seamlessly in diverse environments and lighting conditions. Early methods (von Marcard et al., 2017) used six IMUs for an optimization-based offline pose reconstruction, while later advancements (Huang et al., 2018; Yi et al., 2021) employed deep learning-based models for real-time accuracy. A key challenge with IMUs is their lack of positional data, leading to inaccuracies in global position estimation. Recent studies, such as Jiang et al. (2022b), have introduced Transformer-based models to address this issue, whereas Yi et al. (2022) employs physics-based approaches to achieve more realistic motion. However, IMU methods frequently encounter root positional drift over time (Ami-Williams et al., 2023). In efforts to mitigate drift, researchers such as Guzov et al. (2021); Yi et al. (2023); Lee and Joo (2024) have integrated monocular cameras with SLAM algorithms, aiming for more precise localization.
Motion Capture with Egocentric Cameras
Pose estimation using egocentric cameras has garnered considerable attention in recent years, mainly due to offering a markerless setup. However, there are common challenges inherent to computer vision, such as occlusion and varying lighting conditions. Addressing these challenges, Shiratori et al. (2011) and Rhodin et al. (2016) were among the first to employ body-worn and dual fisheye cameras for motion capture. The technique was further refined by Xu et al. (2019), demonstrating the efficacy of a single head-mounted fisheye camera. Due to recent advancements in deep learning models, fisheye camera-based pose estimation enhancements were subsequently made (Jiang and Ithapu, 2021; Tome et al., 2019; Wang et al., 2021; Tome et al., 2023; Wang et al., 2023). Zhang et al. (2021) tackled the issue of lens distortion through automatic calibration, whereas the works of Yuan and Kitani (2018, 2019) incorporated physics-based controllers to achieve realistic poses. In a more recent development, Li et al. (2023) introduces a two-stage estimation process, presenting a method that simplifies motion capture without relying on paired egocentric videos and human motions. Despite significant advancements in the field, occlusions remain a persistent challenge in computer vision techniques, often necessitating the integration of additional sensor data to ensure robust pose predictions. Moreover, the incorporation of cameras in consumer-grade systems raises concerns regarding user privacy, highlighting the need for careful consideration in the design and application of these technologies.
@@
The increasing accessibility of commercial VR devices has led to significant progress in reconstructing full-body poses using 6-DoF trackers, which capture both position and rotation. These trackers, utilizing external sensors or computer vision, ensure accurate global information. Early work by Dittadi et al. (2021) used a variational autoencoder (Kingma and Welling, 2013) for pose reconstruction from three tracking points, though without global translation estimates. Subsequent studies by Ahuja et al. (2021) and Ponton et al. (2022) explored matching user poses to a motion dataset, similar to the approach in Motion Matching (Clavet, 2016). Adding another tracker on the pelvis, Yang et al. (2021) developed a model using Gated Recurrent Units (GRUs) to predict lower-body movements with velocity data. To capture the continuous nature of motion, Jiang et al. (2022a) and Zheng et al. (2023) employed Transformer encoders for real-time pose estimation. Similarly, various generative AI-based methods have been proposed, including conditional flow-based models (Aliakbarian et al., 2022) and diffusion models (Du et al., 2023; Castillo et al., 2023). Autoencoders have played a pivotal role in developing structured latent spaces for pose reconstruction. In this context, Shin et al. (2023) focused on navigating the latent space to reconstruct poses from sparse input, while Milef et al. (2023) introduced new interpolation and pose anomaly detection methods to avoid traversing invalid regions of the latent space. Additionally, Ponton et al. (2023) combined a skeleton-aware autoencoder with learned inverse kinematics for precise full-body pose reconstruction, ensuring better accuracy in end-effectors. Other techniques involve the development of reinforcement learning frameworks to create natural and realistic movements, as proposed by Winkler et al. (2022) and Ye et al. (2022), with Lee et al. (2023) extending this to include environment interactions. These techniques have demonstrated their ability to reconstruct high-quality, smooth motion. However, they are often constrained by the rigidity of their deep-learning architectures, which limits their adaptability to varying numbers of sparse inputs or constraints. Furthermore, these methods frequently face challenges in accurately representing hard constraints, such as precise end-effector positioning.
2.2. Learned Inverse Kinematics
@@
In robotics and computer animation, IK solvers are essential for determining the positions and orientations of intermediate joints in a kinematic chain, given the target positions and orientations of end-effectors. Aristidou et al. (2018) offers a comprehensive review of popular IK approaches for human motion reconstruction. Yet, traditional IK solvers often face scalability issues for multi-chain characters and a balance must be struck between computational efficiency and pose naturalness (Caserman et al., 2019).
In robotics, learned IK methods utilize neural networks to solve or explore solutions for target end-effectors. These networks range from lightweight feed-forward networks (Bócsi et al., 2011; Duka, 2014; Csiszar et al., 2017; Bensadoun et al., 2022) to more complex structures like generative adversarial networks (Ren and Ben-Tzvi, 2020) and conditional normalizing flow networks (Ames et al., 2022). While they speed up IK computations, they have only been studied for limited kinematic chains, such as robotic arms with fewer degrees of freedom than the human body.
@@
In computer animation, early work by Grochow et al. (2004) and Wu et al. (2011) introduced a data-driven IK system using Gaussian processes for versatile pose modeling. The work by Huang et al. (2017) further enhanced the idea of using Gaussian models with a traditional Jacobian-based solver for real-time pose generation. With the augmented capabilities of modern deep-learning-based architectures, Victor et al. (2021) proposes an autoencoder-based IK solver. A common issue in deep learning-based IK methods is their fixed network structure, which limits their use to a specific number of sparse inputs. To address this, Oreshkin et al. (2021) implemented prototype encoding and residual connections (ProtoRes), which enables dynamic adjustments on the number of end-effectors. Building on this, Voleti et al. (2022) adapted the method for SMPL models (Loper et al., 2015) and incorporated initial pose estimation from images. Further advancements were made by Agrawal et al. (2023), who refined the technique to ensure pose consistency across changes. They utilized a skeletal graph structure, enabling the encoding of hard constraints by restricting information flow at certain joints. This adaptation significantly enhances the workflow for artists by allowing them to maintain base poses while editing specific body parts. However, protores-based systems still contain predefined constraints embedded within the network structure, limiting the ability to modify these constraints dynamically. Moreover, as both constraints and end-effectors variability are encoded as inputs, the network requires training across all possible permutations of sparse input and constraints, posing scalability challenges.
3. Background
@@
(2) $-D_{KL}\left(q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x}^{(i)})\,||\,p(\mathbf{z})\right)+\mathbb{E}_{\mathbf{z}\sim q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x}^{(i)})}\left[\log p_{\boldsymbol{\theta}}(\mathbf{x}^{(i)}|\mathbf{z})\right]$
@@
(3) $q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x}^{(i)})=\mathcal{N}\left(\mu_{\boldsymbol{\phi}}(\mathbf{x}^{(i)}),\mathbf{I}\,\sigma_{\boldsymbol{\phi}}(\mathbf{x}^{(i)})\right)$

### v2

@@ Abstract
High-quality motion reconstruction that follows the user’s movements can be achieved by high-end mocap systems with many sensors. However, obtaining such animation quality with fewer input devices is gaining popularity as it brings mocap closer to the general public. The main challenges include the loss of end-effector accuracy in learning-based approaches, or the lack of naturalness and smoothness in IK-based solutions. In addition, such systems are often finely tuned to a specific number of trackers and are highly sensitive to missing data, e.g., in scenarios where a sensor is occluded or malfunctions. In response to these challenges, we introduce DragPoser, a novel deep-learning-based motion reconstruction system that accurately represents hard and dynamic constraints, attaining real-time high end-effectors position accuracy. This is achieved through a pose optimization process within a structured latent space. Our system requires only one-time training on a large human motion dataset, and then constraints can be dynamically defined as losses, while the pose is iteratively refined by computing the gradients of these losses within the latent space. To further enhance our approach, we incorporate a Temporal Predictor network, which employs a Transformer architecture to directly encode temporality within the latent space. This network ensures the pose optimization is confined to the manifold of valid poses and also leverages past pose data to predict temporally coherent poses. Results demonstrate that DragPoser surpasses both IK-based and the latest data-driven methods in achieving precise end-effector positioning, while it produces natural poses and temporally coherent motion. In addition, our system showcases robustness against on-the-fly constraint modifications, and exhibits adaptability to various input configurations and changes. The complete source code, trained model, animation databases, and supplementary material used in this paper can be found at https://upc-virvig.github.io/DragPoser
{CCSXML}
<ccs2012> <concept> <concept_id>10010147.10010371.10010352.10010238</concept_id> <concept_desc>Computing methodologies Motion capture</concept_desc> <concept_significance>500</concept_significance> </concept> <concept> <concept_id>10010147.10010371.10010352.10010380</concept_id> <concept_desc>Computing methodologies Motion processing</concept_desc> <concept_significance>500</concept_significance> </concept> <concept> <concept_id>10010147.10010371.10010352</concept_id> <concept_desc>Computing methodologies Animation</concept_desc> <concept_significance>500</concept_significance> </concept> <concept> <concept_id>10010147.10010257.10010258</concept_id> <concept_desc>Computing methodologies Learning paradigms</concept_desc> <concept_significance>300</concept_significance> </concept> </ccs2012>
[…]
volume: 44issue: 2
1 Introduction
@@
Human motion reconstruction and editing have gained significant attention in diverse industries such as entertainment, sports, and rehabilitation, with applications spanning virtual reality (VR), augmented reality (AR), training, education, simulations, and collaborative work. In professional settings, motion reconstruction typically relies on high-quality optical motion capture (mocap) systems or Inertial Measurement Unit (IMU)-based systems equipped with numerous sensors. Although effective, these systems come with significant costs and require extensive calibrations and technical expertise. Consequently, recent trends indicate a growing demand for cost-effective and minimally intrusive mocap systems that leverage consumer-grade hardware for motion reconstruction. This shift in focus aims to cater to applications in the Metaverse, targeting affordability and accessibility. Motion editing often involves the use of non-learned inverse kinematics (IK) methods such as CCD [kenwright:2012] or FABRIK [aristidou:2011]. While IK enables the rapid creation of animations using end-effectors, it frequently leads to time-incoherent or unnatural human motion due to the inherently underdetermined nature of the problem.
@@
With the growing availability of extensive repositories of motion capture data and the rapid advancements in deep learning, various data-driven methodologies have emerged to grapple with the intricacies and challenges of human motion reconstruction. Recent studies have focused on the reconstruction of poses from sparse input [ponton2023sparseposer, yi:2023, Winkler:2022], aiming to reduce reliance on expensive motion capture suits. These methodologies, leveraging their learning from vast and precise mocap datasets, reconstruct human motion that is characterized by both temporal coherence and naturalness, overcoming the shortcomings of traditional IK [oreshkin2021a, Agrawal:2023]. However, they often encounter limitations arising from their rigid neural network architectures, restricting users to a specific set of sparse inputs or constraints. Additionally, they are sensitive to faulty sensors or tracking losses. Despite their capacity to reconstruct smooth animations, deep learning-based methods commonly face challenges in dealing with hard constraints, often requiring the fallback to traditional IK methods for refining final poses [Jiang:2022b, Ponton:2022b], which can inadvertently compromise the naturalness of the data-driven results.
@@
This paper introduces DragPoser, an innovative deep-learning-based system designed for motion reconstruction. Unlike conventional methods that rely on direct full-body pose prediction from sparse inputs or optimize short motion sequences within a latent space, DragPoser leverages a structured latent space and employs a pose optimization process to represent both hard and dynamic constraints accurately. The key difference lies in utilizing an optimization process within the learned latent space, enabling dynamic guidance in the pose search. The system undergoes a singular training phase on an extensive human motion dataset. Subsequently, constraints are dynamically defined as losses, and the pose is iteratively refined by computing gradients within the latent space. To ensure temporal coherence and constrain pose optimization within the valid pose manifold, we integrate a Temporal Predictor network, which directly encodes temporality into the latent space. This departure from traditional direct pose prediction methods enables a more precise and versatile exploration of poses, providing increased generality and improved accuracy when enforcing rigid constraints on end effector positions.
@@
We extensively tested and compared our method against state-of-the-art methods in a variety of challenging motion sequences from public datasets, with different sensor configurations. Our results show that DragPoser outperforms both traditional IK-based methods and the newest data-driven approaches in nearly all error metrics. DragPoser excels at providing accurate end-effector positioning (as in traditional IK systems) but at the same time, it generates natural poses and temporally coherent motion, a feat that data-driven methods typically achieve only in the absence of hard constraints. DragPoser, however, boasts another distinct advantage: its adaptability to hardware changes. We showcase this adaptability by demonstrating DragPoser’s robust capability to effectively handle missing input data caused by occlusions, faulty sensors, or communication problems. Additionally, users can dynamically add constraints in real-time to target specific types of motions. This dual adaptability not only reinforces the system’s resilience but also enhances its performance in real-world applications.
2 Related Work
@@
2.1 Motion Capture with Sparse Input
@@
Recent research in human motion reconstruction has focused on utilizing fewer Inertial Measurement Units (IMUs) attached to the body. This approach eliminates the need for external sensors or cameras, requires no line-of-sight, and operates seamlessly in diverse environments and lighting conditions. Early methods [Marcard:2017] used six IMUs for an optimization-based offline pose reconstruction, while later advancements [Huang:2018, Yi:2021] employed deep learning-based models for real-time accuracy. A key challenge with IMUs is their lack of positional data, leading to inaccuracies in global position estimation. Recent studies [Jiang:2022], have introduced Transformer-based models to address this issue, whereas Yi et al. [Yi:2022] employs physics-based approaches to achieve more realistic motion. However, IMU methods frequently encounter root positional drift over time [Ami-Williams:2023]. In efforts to mitigate drift, researchers [guzov:2021, yi:2023, lee:2024] have integrated monocular cameras with SLAM algorithms, aiming for more precise localization.
@@
The increasing accessibility of commercial VR devices has led to significant progress in reconstructing full-body poses using 6-DoF trackers, which capture both position and rotation. These trackers, utilizing external sensors or computer vision, ensure accurate global information. Early work [Dittadi:2021] used a variational autoencoder for pose reconstruction from three tracking points, though without global translation estimates. Subsequent studies [Ahuja:2021, Ponton:2022b] explored matching user poses to a motion dataset, similar to the approach in Motion Matching [Clavet:2016]. Adding another tracker on the pelvis, Yang et al. [Yang:2021] developed a model using Gated Recurrent Units (GRUs) to predict lower-body movements with velocity data. To capture the continuous nature of motion, Jiang et al. [Jiang:2022b] and Zheng et al. [zheng2023] employed Transformer encoders for real-time pose estimation. Similarly, various generative AI-based methods have been proposed, including conditional flow-based models [Aliakbarian:2022] and diffusion models [du2023a, castillo2023]. Autoencoders have played a pivotal role in developing structured latent spaces for pose reconstruction. In this context, Milef et al. [milef2023] introduced new interpolation and pose anomaly detection methods to avoid traversing invalid regions of the latent space. Additionally, Ponton et al. [ponton2023sparseposer] combined a skeleton-aware autoencoder with learned inverse kinematics for precise full-body pose reconstruction, ensuring better accuracy in end-effectors. Other techniques [Winkler:2022, ye2022] involved the development of reinforcement learning frameworks to create natural and realistic movements. Lee et al. [lee2023] extended this to include environment interactions. These techniques have demonstrated their ability to reconstruct high-quality, smooth motion. However, they are often constrained by the rigidity of their deep-learning architectures, which limits their adaptability to varying numbers of sparse inputs or constraints. Furthermore, these methods frequently face challenges in accurately representing hard constraints, such as precise end-effector positioning.
Motion priors for Motion Capture
Leveraging compressed latent space representations has become essential for human motion synthesis due to the challenges of working directly with large amounts of mocap data. Early work [andrews:2016] used a physics-based model and a motion prior to reconstruct poses from a set of IMUs and optical markers. Subsequent work [ling:2020, peng:2022] explored VAEs for encoding and decoding poses, thereby enabling deep reinforcement learning in the latent space. In the context of learning neural representations as priors, Liu et al. [liu_learning_2022] introduced Lipschitz regularization to enforce continuity and facilitate smooth interpolation and extrapolation operations. Motion priors have proven valuable for handling noisy or faulty pose data. For instance, Rempe et al. [rempe:2021] achieved impressive results by training a motion prior on the AMASS dataset, later refined by Shi et al. [shi:2023] with a periodic autoencoder [starke:2022]. Once their latent space is trained, pose transition sequences are optimized to satisfy some given constraints, such as end-effector positions or fitting a point cloud. However, challenges remain in reproducing high-frequency details and real-time processing. Working with transitions limits the ability to make targeted adjustments to individual poses within the latent space, as each pose is dependent on the entire preceding sequence. Our approach addresses these limitations by representing full poses directly in the latent space. Direct pose optimization in the latent space allows our method to have a more fine-grained control while being performant, and ensures constraints are met frame by frame instead of optimizing large motion sequences.
2.2 Learned Inverse Kinematics
@@
In computer animation, IK solvers are essential for determining the positions and orientations of intermediate joints in a kinematic chain, given the target positions and orientations of end-effectors. Aristidou et al. [Aristidou:2018] offers a comprehensive review of popular IK approaches for human motion reconstruction. Yet, traditional IK solvers often face scalability issues for multi-chain characters and a balance must be struck between computational efficiency and pose naturalness [Caserman:2019].
@@
Early work [Grochow:2004, Wu:2011] introduced a data-driven IK system using Gaussian processes for versatile pose modeling. The work by Huang et al. [Huang:2017] further enhanced the idea of using Gaussian models with a traditional Jacobian-based solver for real-time pose generation. With the augmented capabilities of modern deep-learning-based architectures, Victor et al. [Victor:2021] proposed an autoencoder-based IK solver. A common issue in deep learning-based IK methods is their fixed network structure, which limits their use to a specific number of sparse inputs. To address this, Oreshkin et al. [oreshkin2021a] implemented prototype encoding and residual connections (ProtoRes), which enables dynamic adjustments on the number of end-effectors. Building on this, Voleti et al. [voleti2022] adapted the method for SMPL models [loper2015] and incorporated initial pose estimation from images. Further advancements were made by Agrawal et al. [Agrawal:2023], who refined the technique to ensure pose consistency across changes. They utilized a skeletal graph structure, enabling the encoding of hard constraints by restricting information flow at certain joints. This adaptation significantly enhances the workflow for artists by allowing them to maintain base poses while editing specific body parts. However, protores-based systems still contain predefined constraints embedded within the network structure, limiting the ability to modify these constraints dynamically. Moreover, as both constraints and end-effectors variability are encoded as inputs, the network requires training across all possible permutations of sparse input and constraints, posing scalability challenges.
3 Background
@@
$-D_{KL}\left(q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x}^{(i)})\,||\,p(\mathbf{z})\right)+\mathbb{E}_{\mathbf{z}\sim q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x}^{(i)})}\left[\log p_{\boldsymbol{\theta}}(\mathbf{x}^{(i)}|\mathbf{z})\right]$ (2)
@@
$q_{\boldsymbol{\phi}}(\mathbf{z}|\mathbf{x}^{(i)})=\mathcal{N}\left(\mu_{\boldsymbol{\phi}}(\mathbf{x}^{(i)}),\mathbf{I}\,\sigma_{\boldsymbol{\phi}}(\mathbf{x}^{(i)})\right)$ (3)